# PartComposer: Learning and Composing Part-Level Concepts from Single-Image Examples

### v1

@@ Appendix B Comparison Implementation
Break-a-Scene (Avrahami et al., 2023) is originally designed to learn mutiple subject-level concepts in a single image. We modify the data loading and processing scheme to support learning part-level concepts from single-image examples. In addition, since we can only access 24G VRAM GPU (RTX3090), we add LoRA (Hu et al., 2022) following standard StableDiffusion v2.1 model finetuning scheme to train the LoRA weights instead of the entire diffusion U-Net. This approach has been validated by PartCraft (Ng et al., 2024) that it will not harm the concept learning quality when compared to finetuning the entire U-Net.
@@
PartCraft (Ng et al., 2024) is originally designed to learn part-level concepts from a large dataset of images. Specifically, they require to have 10-20 images for a subject (e.g., a specific bird species.). We modify the data loading and processing scheme to load single-image examples dataset into their pipeline. In addition, we directly provide the part-level masks instead of using their automatic concept discovering method to ensure fair comparison. We use StableDiffusion v1.5 as described in their paper to learn the concepts. Their approach show unstable performance. The results we show in Figure 5 are already the best results we observed. The image quality degrades quickly when the training step number increases (after the step we shown in Figure 5). We also tried to use smaller learning rate but their pipeline struggle to even learn any meaningful concepts. We hypothesis the reason for this is that their method is designed to use on a large dataset instead of the single-image examples in our case.
@@
MuDI (Jang et al., 2024) is originally designed to learn and disentangle similar subject-level concepts from multi-image examples. They generally require over 5 images per subject. We use the SDXL with LoRA as described in their paper to finetune the diffusion model. For the inference, we use their official inference code which first creates a mean-shifted noise by pasting intended part onto a blank image and then add noise, and then run SDXL (Podell et al., 2023) to iteratively denoise. The results we show in Figure 5 is already the best results we observed. Their approach struggle at part-level and with more than 5 concepts - the single-image example of 2 chairs contains 8 concepts.
@@
PiT (Jang et al., 2024) is originally designed for creative visual designs by training priors for different object categories and infer creative designs given different part inputs. This method only require test-time inference, which is different with all the above methods and our methods which are optimization-based. Since we don’t have access to their full training data, we directly use their released 3 checkpoints (product, creature, and toy) to conduct evaluation. To enable fair comparison, we first use 2 ”creature” images that are used in their training and inference through their released ”creature” checkpoint. The goal is to evaluate their concept learning and composition capability for in-distribution data. We then use 2 ”chair” images and inference through their ”product” checkpoint, which is the most related checkpoint in terms of category similarity. The goal for this design is to evaluate their capability to encode and compose parts for slightly out-of-distribution examples.
@@ C.1. Intra-category Results
We conduct experiments on various object categories to demonstrate our model’s capability in learning and re-composing part-level concepts. In each case, objects are decomposed into over four semantic parts, and our method learns to mix these parts across instances from the same category. We want to note that all the semantic part naming are annotated just for ease-of-read in the paper. We do not require semantically mearningful part decomposition and detailed annotation for the part-level concepts in our pipeline since we only treat them as visual information.
@@
Parts are: armrest, backrest, legs, and seat. Figure 11 and Figure 12 show concept composition results across 6 different chair pairs.
@@
Parts are: headboard, base, mattress, and pillow. Figure 13 shows concept composition for a pair of beds.
@@
Parts are: base, stand, seat, and weight. Figure 14 shows concept composition for a pair of gym equipment examples.
@@
Parts are: front, cockpit, tail, and wheels. Figure 15 shows concept composition for three pairs of vehicles.
@@
Parts are: handle, frame, seat, and wheels. Figure 16 shows concept composition results for a pair of bikes.
@@
Parts are: body, wing, tail, and engine. Figure 17 shows concept composition results for a pair of planes.
@@
Parts are: head, body, arm, and leg. Figure 18 shows concept composition results for a pair of robots.
@@
Parts are: head, body, wings, tail, and legs. Figure 19 shows concept composition for a bird and a creature pair.
@@
Figure 20 shows concept composition for a mushroom-like character and a santa-like character. Parts are: head, eyes, face, body, and legs. Figure 20 shows concept composition for a hermit-like character and a reptile-like character. This example aim to demonstrate that our method can learn a large number (i.e., in this case, 7 parts per image) of fine-grained concepts for parts and re-compose them. Parts are: head, eyes, face, nose, arms, body, and legs.
@@ C.2. Cross-category Results
To demonstrate generalization beyond intra-category recomposition, we evaluate our method on cross-category part composition. Figure 21 shows 4 hybrid compositions: a chair and a wheel chair, a gym equipement and a wheel chair, a chair and a bed, and a chair and a gym equipment..
@@
Appendix D Further Analysis
Visualization for concept predictor
We visualize the convergence of our concept predictor’s segmentation head to demonstrate the effectiveness of maximizing mutual information. Figure 22 illustrates the comparison of segmentation results from concept predictor and the ground truth masks. The concept predictor gradually converge to predict correct segmentation of the parts, showing the effectiveness our concept predictor design.
Figure 22. Visualization for the convergence of segmentation head in our concept predictor.
Weight and Scale Experiments
The most important weight in our experiments is the weight for the concept predictor loss $\lambda_{\text{info}}$ explained in Section 3. We found that $\lambda_{\text{info}}=0.05$ is usually the best weight for most examples, though some examples require slight higher weight ranging from $0.075$ to $0.1$. Increasing the $\lambda_{\text{info}}$ might show obvious improvement on concept preserving but may result in un-natural artifacts since the diffusion prior may get damaged. Figure 23 illustrates this phenomena. When increasing the weight from $0.01$ to $0.05$, concepts identity preservation is improved, showing the effectiveness of our concept predictor. When further increasing the weight, although concepts are still well-preserved, there are some visual artifacts including saturation shift.
Figure 23. Visualization of the impacts of different information loss weight
Appendix E Limitations
Although our method can learn well entangled part-level concepts can re-compose them, the image generation quality for some challenging inputs might still need improvement. For example, when composing parts from a truck and s sports car in Figure 15, there are some noticable artifacts in the generated images even though all the intended part concepts are preserved. Our pipeline also sometimes do not preserve the exact details for very thing or high-frequency structures like the horizontal bars in chair legs (in Figure 5) and the number of bird legs (in Figure 19). This problem might partially due to the inherent limitation of text-to-image diffusion models since these models can produce images containing unrealistic details like a bird with 3 or more legs. Additionally, for cross-category part re-composition, not all compositions of parts result in meaningful virtual objects. For example, when mixing a chair and a gym equipment in Figure 21, the composition in the third column results in non-meaningful objects. A composition prediction scheme may be developed in this case to predict the possible meaningful part compositions instead of naively trying all possible compositions.

### v2

@@ Appendix B Comparison Implementation
Break-a-Scene (Avrahami et al., 2023) is originally designed to learn multiple subject-level concepts in a single image. We modify the data loading and processing scheme to support learning part-level concepts from single-image examples. In addition, since we can only access 24G VRAM GPU (RTX3090), we add LoRA (Hu et al., 2022) following the standard StableDiffusion v2.1 model finetuning scheme to train the LoRA weights instead of the entire diffusion U-Net. This approach has been validated by PartCraft (Ng et al., 2024) that it will not harm the concept learning quality when compared to finetuning the entire U-Net.
@@
PartCraft (Ng et al., 2024) is originally designed to learn part-level concepts from a large dataset of images. Specifically, they require to have 10-20 images for a subject (e.g., a specific bird species). We modify the data loading and processing scheme to load the single-image examples dataset into their pipeline. In addition, we directly provide the part-level masks instead of using their automatic concept discovery method to ensure fair comparison. We use StableDiffusion v1.5 as described in their paper to learn the concepts. Their approach shows unstable performance. The results we show in Fig. 5 of the main paper are already the best results we observed. The image quality degrades quickly when the training step number increases (after the step we showed in Fig. 5 of the main paper). We also tried to use a smaller learning rate, but their pipeline struggled to even learn any meaningful concepts. We hypothesize that the reason for this is that their method is designed to be used on a large dataset instead of the single-image examples in our case.
@@
MuDI (Jang et al., 2024) is originally designed to learn and disentangle similar subject-level concepts from multi-image examples. They generally require over 5 images per subject. We use the SDXL with LoRA as described in their paper to finetune the diffusion model. For the inference, we use their official inference code, which first creates a mean-shifted noise by pasting the intended part onto a blank image, and then adds noise, and then run SDXL (Podell et al., 2023) to iteratively denoise. The results we show in Fig. 5 of the main paper are already the best results we observed. Their approach struggles at the part-level and with more than 5 concepts - the single-image example of 2 chairs contains 8 concepts.
@@
PiT (Jang et al., 2024) is originally designed for creative visual designs by training priors for different object categories and inferring creative designs given different part inputs. This method only requires test-time inference, which is different from all the above methods and our methods, which are optimization-based. Since we don’t have access to their full training data, we directly use their released 3 checkpoints (product, creature, and toy) to conduct evaluation. To enable fair comparison, we first use 2 ”creature” images that are used in their training and inference through their released ”creature” checkpoint. The goal is to evaluate their concept learning and composition capability for in-distribution data. We then use 2 ”chair” images and inference through their ”product” checkpoint, which is the most related checkpoint in terms of category similarity. The goal for this design is to evaluate their capability to encode and compose parts for slightly out-of-distribution examples.
@@ C.1. Intra-category Results
We conduct experiments on various object categories to demonstrate our model’s capability in learning and re-composing part-level concepts. In each case, objects are decomposed into over four semantic parts, and our method learns to mix these parts across instances from the same category. We want to note that all the semantic part names are annotated just for ease of reading in the paper. We do not require semantically meaningful part decomposition and detailed annotation for the part-level concepts in our pipeline since we only treat them as visual information.
@@
Parts are: armrest, backrest, legs, and seat. Fig. 13 and Fig. 14 show concept composition results across 6 different chair pairs.
@@
Parts are: headboard, base, mattress, and pillow. Fig. 15 shows concept composition for a pair of beds.
@@
Parts are: base, stand, seat, and weight. Fig. 16 shows concept composition for a pair of gym equipment examples.
@@
Parts are: front, cockpit, tail, and wheels. Fig. 17 shows concept composition for three pairs of vehicles.
@@
Parts are: handle, frame, seat, and wheels. Fig. 18 shows concept composition results for a pair of bikes.
@@
Parts are: body, wing, tail, and engine. Fig. 19 shows concept composition results for a pair of planes.
@@
Parts are: head, body, arm, and leg. Fig. 20 shows concept composition results for a pair of robots.
@@
Parts are: head, body, wings, tail, and legs. Fig. 21 shows concept composition for a bird and a creature pair.
@@
Fig. 22 shows concept composition for a mushroom-like character and a santa-like character. Parts are: head, eyes, face, body, and legs. Fig. 22 shows concept composition for a hermit-like character and a reptile-like character. This example aim to demonstrate that our method can learn a large number (i.e., in this case, 7 parts per image) of fine-grained concepts for parts and re-compose them. Parts are: head, eyes, face, nose, arms, body, and legs.
@@ C.2. Cross-category Results
To demonstrate generalization beyond intra-category recomposition, we evaluate our method on cross-category part composition. Fig. 23 shows 4 hybrid compositions: a chair and a wheel chair, a gym equipement and a wheel chair, a chair and a bed, and a chair and a gym equipment..